OpenStudy (anonymous):

Find the value(s) of c such that the area of the region bounded by the parabolae y=x^2−c^2 and y=c^2−x^2 is 4608.

OpenStudy (amistre64):

whats half the area?

OpenStudy (amistre64):

or for that matter, a quarter of it?

OpenStudy (amistre64):

\[\int_{0}^{b}c^2-x^2~dx=\frac{4608}{4}\]

OpenStudy (amistre64):

looks like 0 to c is more meaningful to us since: x^2 - c^2 = c^2 - x^2 when c=x

OpenStudy (anonymous):

How do you find b?

OpenStudy (triciaal):